By using Slopes form, show that (3, 1), (6, 3) and (2, 9) are the vertices of a right triangle.
Answer by dolly(163)

hi,
The slope of the line joining (3,1) and (6,3) = (3-1)/(6-3)
= 2/3 -------------(1)
The slope of the line joining (6,3)and (2,9) = (9-3)/(2-6)
= 6/-4
= 3/-2 ------------(2)
Multiplying (1) and (2) we get, (2/3)(3/-2)
= -1
As the product of the slopes is -1, the lines are perpendicular .
==> the lines intersect at right angle
==> the given points are the vertices of a right triangle
good luck!!!
dolly